OpenStudy (lgbasallote):

LGBARIDDLE 2 + 7 = 16 3 + 5 = 21 4 + 4 = 28 5 + 2 = 30 6 + 8 = ???

OpenStudy (ajprincess):

6*5+8*6=30+48=78

hartnn (hartnn):

2*7 + 2*1 = 16 3*5 + 3*2 = 21 4*4 + 4*3 = 28 so 6*8 + 6*5 = 78 as correctly done by @ajprincess

OpenStudy (lgbasallote):

clever @ajprincess

OpenStudy (ajprincess):

Thanx